OpenStudy (kaiz122):

a tablet is 7m high is placed on a wall with its base9 m above the level of an observer's eye. how far from the wall should an observer stand in order that the angle subtended at the eye by the table be a maximum?

OpenStudy (kaiz122):

help me please @saifoo.khan @ParthKohli @KatrinaKaif

OpenStudy (shubhamsrg):

you attempting/preparing for ISI ? coz this is a ques of ISI sample paper..

OpenStudy (shubhamsrg):

 we have to find x for maximum y now, tan z = 9/x tan(y+z) = 16/x =>(tan y + tanz) /(1- tany tanz) = 16/x =>tan y + 9/x = (16/x) (1- tany 9/x ) => tan y +9/x = 16/x - 144/x^2 tany tan y = (7/x) / (1+ 144/x^2) = 7x/(144 + x^2) = 7 / (144/x +x) now for y maximum,,tany will be maximum,,since y is acute. now tany will be maximum for denominator of RHS be minimum i.e. our problem is to minimize (144/x + x) putting d/dx( 144/x +x) =0 ; => - 144/x^2 + 1 =0 => x= 12 ,,which should be our ans you can confirm this is the minimum value as d2 /dx^2 of this is +ve. hope i was clear enough..